Pearl Corporation bought a machine on June 1, 2015, for $33,170, f.o.b. the place of manufacture. Freight to the point where it was set up was $214, and $535 was expended to install it. The machine’s useful life was estimated at 10 years, with a salvage value of $2,675. On June 1, 2016, an essential part of the machine is replaced, at a cost of $2,119, with one designed to reduce the cost of operating the machine. The cost of the old part and related depreciation cannot be determined with any accuracy.

On June 1, 2019, the company buys a new machine of greater capacity for $37,450, delivered, trading in the old machine which has a fair value and trade-in allowance of $21,400. To prepare the old machine for removal from the plant cost $80, and expenditures to install the new one were $1,605. It is estimated that the new machine has a useful life of 10 years, with a salvage value of $4,280 at the end of that time. (The exchange has commercial substance.)

Assuming that depreciation is to be computed on the straight-line basis, compute the annual depreciation on the new equipment that should be provided for the fiscal year beginning June 1, 2019. (Round answer to 0 decimal places, e.g. 45,892.)

Depreciation for the year beginning June 1, 2019
Answer #1

The cost of assets include costs for bringing it to use i.e. freight and installation to getting it ready for use.

When the business exchanges fixed assets with another and transaction has commercial substance, the business records the asset acquired at its fair value or fair value of assets given up, whichever is more readily available. When no commercial substance exists, the asset swap has no accounting impact as there is no significant change. However, this may result in profit and loss.

June 1, 2015

Purchase

$ 33,160

Add: Freight

214

Add: Installation

       535

Total cost

$ 33,909

           Annual depreciation charge: ($33,909 – $2,675) ÷ 10 = $3,123

           On June 1,2016, the essential part of machine is replaced at $2119. This shall be added to the cost of machine as 33,909+2119 = $ 36,028 which will be revised cost.

           Revised depreciation would be (36,028-3123-2675)/9 = 30230/9 = 3359

           The machine is used for 3 years until latter swapped with asset which transaction has commercial substance.

Book value, old machine, June 1, 2019:

   [$36,028 – $3,123 – ($3,359 X 3)]

$ 22,828

Fair value

(21,400)

Loss on exchange

1428

Cost of removal

80

Total loss

1508

New Machine

Basis of new machine

Cash paid ($37,450 – $21,400)

16,050

Fair value of old machine

21,400

Installation cost

  1,605

Total cost of new machine

$ 39,055

Depreciation for the year beginning June 1, 2019 = ($39,055 – $4,280) ÷ 10 = $ 3478.

kindly upvote